Preliminary Agenda for the VISTA IOT on Nov xx, 2011 ==================================================== (15-16 Garching, 14-15 UK, 9-10 Chile) 1. VIRCAM/VISTA general technical status and planned activities (TSz, VDI) 2. Summary of the Report on the VISTA public surveys 1.5 year into operations (MAr) 3. Survey progres rate, correcting measures for improvement (MRe, MHi, TSz, VDI, etc) 4. Possible data quality improvements (VDI) 5. Software (SADT/OT/P2PP/etc) issues (MRe, MHi, JEm) 6. QC policy change (WHu), data reduciton issues (MIr) 7. Action Items review (enclosed bellow is the open AIs from the last IOT) 8. Other issues 9. Date of the next meeting, proposed May 2012 ************************************************************ ====================== 5. Action Items review ====================== OLD AIs: -------- AI 2009-11: TSz should find out if the constraints sets could be carried with the OBs, to the OS and the fits header of the VIRCAM files. Otherwise it is not easy to propagate these data by interfaces. 2009-07-29: TSz iterating with R. Schmutzer, but DFI (T. Bierwirth) will also need to be involved. For surveys, given the small range of programs it is expected that there are not too many different OBs, and this might need to be followed in a different/manual way in the beginning this could imply a major change for the tools that cannot be done in the last moment. 2010-03-23: the keywords have been implemented except for the moon; there is a ticket with details; some testing done by VDI, no errors found; keep ongoing, try to complete during my next shift end of Mar early Apr. 2010-04-23: VDI reports one formula is missing 2010-05-27: Corresponding PPRS is still pending. VDI asks to keep this AI open until the PPRS is closed. 2010-06-29: No changes with respect to last meeting. VDI asks to keep this AI kept open/pending until the PPRS is closed. 2010-08-24: VDI not present at the meeting. It is not known if there was any progress. Therefore decided to keep this pending while waiting for final status from VDI. 2010-10-05: VDI: still pending, will force the issue with the SW during my next shift from Oct 19 to 30, 2010 STATUS: PENDING ** AI 2010-06: TSz+VDI to improve the TIO training, to prepare an operational manual 2010-05-27: Operations wiki has been updated. More training needed. 2010-06-29: Not much was done due to problems with IQ + intervention + work on IP. Keep the AI open. 2010-08-24: SMI comments that the training is an ongoing task. Therefore it seems that this could be closed. It is not clear about the status of the operational manual mentioned in the AI. Thus for the moment kept as PENDING. Comment from VDI and TSZ on the status required. 2010-10-05: to be kept open for now as a reminder STATUS: PENDING ** AI 2010-09: photometric zeropoints as calculated by the pipeline. 2010-05-27: Ongoing, see corresponding discussion this meeting 2010-06-29: TSZ points out that this is not well defined AI. MIR adds that the currently existing plots already would have shown the problems with zero points from pawprints to tiles. It is not clear whether this is about the zero points from Garching pipeline? WHU states that Garching pipeline has de-blending on. MIR: There is no significant increase in the scatter with respect to what one would expect from the random noise. It is agreed in the end to send one example data set and then the comparison should be done directly between Garching and CASU. 2010-08-24: This is still pending. WHU should send some example to JLE/MIR in order to do the checks and clarify what the problem actually is about. 2010-10-05: the example data were sent to MIr and JLe; the Cambridge pipeline results for the individual detector were as expected from "rms" noise considerations, though offset by 0.1 mag from the Garching results. JLe and WHu are attempting to identify the cause of this difference which is most likely due to differences in calibration frames used. JLe sent a new linearity curve to use; WHu reprocessed the data with the new linearity correction and there were no changes; JLe asked if he could get the calibration frames used from WHu. CASU use the "average" of all detector zero-points to monitor extinction/throughput. STATUS: PENDING ** AI 2011-01: SMi to contact Stefano Zampieri and to find out if the partial transfer of the VISTA data (only calibrations) can be implemented quickly, i.e. on a time-scale of four weeks? 2011-01-27: SMi sent an e-mail already to S.Z. 2011-06-08: done; it is running, the calibration data are transferred on nightly basis STATUS: CLOSED ** AI 2011-02: WHu to compare the ZPs from CASU and Garching from Nov-Dec 2011 to verify the scatter; TSz will acquire more standards during the technical nights in early Apr, after VISTA is back with the recoated mirrors; TSz will be up for the technical time. 2010-06-08 Duplication of AI 2010-09; still pending; CASU got some post-intervention data but the weather was too poor to get good ZP; leave this open until the next IOT in Aug 2011. STATUS: PENDING ** -------- NEW AIs: -------- AI 2011-05: TSz to optimize the AO parameters to get better correction with faint stars; to be discussed with Will Sutherland. STATUS: NEW ** AI 2011-06: MAr et al. to prepare a memo to the management on the VISTA status and survey progress concerning the over-submission in the next period STATUS: NEW ** AI 2011-07: MHi to prepare test OBs with the new SADT; SMi to execute them STATUS: NEW ** AI 2011-08: WHu to send to Klara S. the new ZPs and sky values to update the ETC STATUS: NEW ** AI 2011-09: TSz to investigate if the control loop parameters for the azimuth axis can be modified to speed up the telescope preset. STATUS: NEW ** AI 2011-10: VDI and MRe to investigate if the comments written by TIOs can/should/mast be conveyed to the PIs, especially comment for aborted OBs. This was prompted by user's question. 2011-06-09: the issue was discussed; the comments are relevant only if the aborted OB generated files, if not thee is no need to convey anything but if there were files the comments may help to decide if the data are of any use; decision on this should be postponed until the new night report tool is installed in July; this AI to be kept as a reminder. STATUS: NEW (really, ON HOLD) ** AI 2011-11: JLe to check if the data of the last OB status change is written in the file headers (so the users can see/understand if it is current) STATUS: NEW